# Beyond-Diagonal RIS Architecture Design and Optimization under Physics-Consistent Models

### v1

@@ V-B Optimization of Single-Stream MIMO Systems
For the general physics-consistent channel model given by Proposition 1, $\bar{\mathbf{H}}_{RT}$ has a more complicated expression and is generally nonzero. While a closed-form solution is typically unavailable, we show in the following that the global optimal solution to (19) can be obtained (for fully- and tree-connected RIS) via the semidefinite relaxation (SDR) technique after suitable problem transformations.
To do this, we first substitute the optimal $\mathbf{g}$ into the objective function of (19), which yields
@@
The constant $P_{T}$ does not affect the optimal solution and is thus omitted for brevity. Introducing an auxiliary variable $\mathbf{u}=\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}\mathbf{w}$, problem (20) transforms to
@@
where the non-convex rank-one constraint $\text{rank}(\mathbf{X})=1$ is relaxed. Note that the above problem is defined over the complex space and involves only two constraints. This makes it a tight SDR of (22) [32], that is, there exists a rank-one solution $\mathbf{X}^{*}$ to (24). Therefore, by decomposing $\mathbf{X}^{*}=\mathbf{x}^{*}{\mathbf{x}^{*}}^{H}$, we get the optimal solution to (22).
@@
Solving the ($N_{T}+N_{I}$)-dimensional semidefinite program (SDP) in (24) requires a per-iteration complexity of $\mathcal{O}((N_{T}+N_{I})^{6})$. As the number of RIS elements is typically large, directly solving (24) is computationally expensive. This remark will show that, by carefully exploiting the problem structure, it suffices to solve an SDP with a much lower dimension of $N_{T}+N_{R}$, where $N_{R}\ll N_{I}$ in practice.
@@ V-C Optimization of Multiuser MIMO Systems
Finally, we discuss sum-rate maximization for multiuser MIMO systems. This problem was studied in [10] under the approximate channel model $\mathbf{H}_{\text{app,3}}$, where an alternating direction method of multipliers (ADMM)-based optimization framework applicable to arbitrary RIS architectures was developed. Building on the unified mathematical structure of the general physics-consistent model given in Proposition 1 and $\mathbf{H}_{\text{app,3}}$, the algorithm can be extended to the general physics-consistent model with proper modifications.
@@
For ease of presentation, we assume that each user has a single antenna. In this case, $N_{R}$ is the number of users in the system. The sum-rate maximization problem can be formulated as
[…]
$\|\mathbf{W}\|_{F}^{2}\leq P_{T},$
where $\mathbf{h}_{k}(\bar{\bm{\Theta}})^{H}$ is the $k$-th row of $\mathbf{H}(\bar{\bm{\Theta}})=\bar{\mathbf{H}}_{RT}+\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}$. Similar to [10], we introduce an auxiliary variable $\mathbf{U}=(\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}})^{H}$, then (28) transforms to
@@
where $\mathbf{h}_{k}(\mathbf{U})^{H}$ is the $k$-th row of $\mathbf{H}(\mathbf{U})=\bar{\mathbf{H}}_{RT}+\mathbf{U}^{H}\bar{\mathbf{H}}_{IT}.$ To solve problem (29), we employ an ADMM algorithm similar to that in [10, Eqs. (11a) – (11f)], which first simplifies the sum-rate objective function using fractional programming technique [33, 34] and penalizes the bilinear constraint (29b) into the objective function via the augmented Lagrangian (AL) function, and then alternately updates the variables until convergence. The only difference lies in the update of $(\bar{\mathbf{B}}_{I},\mathbf{B}_{I})$, where the architecture-specific constraint on $\mathbf{B}_{I}$, i.e., $\mathbf{B}_{I}\in\mathcal{B}$, is coupled with $\bar{\mathbf{B}}_{I}$ through the linear constraint (14). In the following, we focus on the update of $(\bar{\mathbf{B}}_{I},\mathbf{B}_{I})$ involved in the ADMM algorithm. Details on the updates for the other variables are omitted for brevity and can be found in [10].
@@ VI Simulation
In this section, we present simulation results to examine the impact of different EM effects and approximations on system performance, using the algorithms proposed in Section V. We focus on fully-connected RIS and evaluate its performance under different channel models. In particular, we investigate two key effects: mutual coupling (MC) among RIS elements and unilateral approximation (UA). To this end, we assume that the source impedances at the transmitter and load impedances at the receiver are equal to the reference impedance $Z_{0}$, set as $Z_{0}=50\,\Omega$, which gives $\mathbf{Z}_{R}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{T}=Z_{0}\mathbf{I}$, and the transmit and receive antennas are perfectly matching with no mutual coupling, which gives $\mathbf{Z}_{TT}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{RR}=Z_{0}\mathbf{I}.$ In addition, the direct link between the transmitter and receiver is assumed to be fully obstructed, i.e., $\mathbf{Z}_{RT}=\mathbf{0}$. We focus on the following three channel models:
@@
We first consider a scenario where the distances between transmitter, RIS, and receiver are large. The elements of $\mathbf{Z}_{RI}$ and $\mathbf{Z}_{IT}$ are generated as independent Gaussian random variables [17], i.e., $\mathbf{Z}_{RI}]_{i,j}\sim\mathcal{CN}(0,\rho_{RI}),$ and $\mathbf{Z}_{IT}]_{i,j}\sim\mathcal{CN}(0,\rho_{IT})$, with path gain $\rho_{IT}=4Z_{0}^{2}10^{-8}$ and $\rho_{RI}=4Z_{0}^{2}10^{-4}$. The transmit power is set as $P_{T}=20$ dBm and the noise power is $\sigma^{2}=-80$ dBm.
@@
To investigate the performance of different channel models for different systems, we solve (18), (19), and (28) under both the general physics-consistent model $\mathbf{H}$ and the approximate models $\mathbf{H}_{\text{app},2}$ and $\mathbf{H}_{\text{app},3}$. In Fig. 2, we depict the receive power for SISO and single-stream MIMO systems and the sum rate for multiuser MISO systems achieved by the three channel models. To better investigate the impact of mutual coupling among RIS antennas, we consider two different values of inter-element spacing, $d=\lambda/2$ and $d=\lambda/4$, respectively. As shown in the figure, the performance with and without the unilateral approximation is almost indistinguishable in all cases. In contrast, mutual coupling among RIS antennas has a noticeable impact on the performance, particularly when the inter-element spacing is small. The presence of mutual coupling improves system performance compared with the case without coupling, and stronger coupling leads to more significant performance gains. We remark here that the benefits of mutual coupling under Rayleigh fading channels were previously reported for SISO systems in [17]. Our results demonstrate that the same conclusion also holds for MIMO and multiuser scenarios.
@@
In Fig. 3, we further evaluate the quality of the approximate solutions obtained using the channel models $\mathbf{H}_{\text{app,2}}$ and $\mathbf{H}_{\text{app,3}}$ to better visualize the accuracy of different approximations. In particular, the quality of each approximate solution is quantified by applying it to the general physics-consistent channel model $\mathbf{H}$ and computing its relative performance compared with the solution obtained from $\mathbf{H}$. From Fig. 3, we can draw similar conclusions as those from Fig. 2. First, the unilateral approximation has a negligible effect in all cases, yielding nearly 100% relative performance compared to the solution of the general physics-consistent model. Second, ignoring mutual coupling among RIS elements leads to performance degradation, particularly when the inter-element spacing is small. For example, when $d=\lambda/4$, the solution from the mutual-coupling-unaware model achieves only around $65\%$ of the performance.
@@
In all the above figures, the distances between transmitter, RIS, and receiver are set to be much larger than the wavelength, which is typically the case in practice. It is commonly accepted that the unilateral approximation accurately represents real world channels under this condition, as also validated by our results. In Fig. 5, we further investigate the range of validity of the unilateral approximation. We consider an extreme scenario where the distance between the transmitter and RIS is on the order of the wavelength. The receiver is still assumed to be located far from the RIS. Specifically, we set the transmitter parallel to the $y$-axis, centered at $(d,0,r\lambda)$, where $d$ is the inter-element spacing of RIS, $\lambda$ is the wavelength, and $r$ is a factor determining the distances between transmitter and RIS. The matrix $\mathbf{Z}_{IT}$ is then computed according to the model in (35), and all other components of the $\mathbf{Z}$ matrix are generated as in the previous simulations. We assume that MC is aware and focus on the effect of unilateral approximation by comparing the performance achieved with models $\mathbf{H}_{\text{app,2}}$ and $\mathbf{H}$. The number of RIS elements is fixed as $N_{I}=64$.
As shown in Fig. 5, the unilateral approximation remains accurate across a broad range, even when the distance between the transmitter and RIS is on the order of the wavelength. For example, even at $0.1\lambda$, the solution obtained under the unilateral approximation still achieves over 95% of the performance of the exact solution for the SISO system. Although the accuracy decreases with the the number of transmit/receive antennas and users, it is still sufficiently high for practical scenarios.
In summary, we can draw the following conclusions from the simulation results regarding the effects of mutual coupling and unilateral approximation. First, stronger mutual coupling among RIS elements can enhance system performance for all SISO, MIMO, and multiuser MISO systems under Rayleigh fading channels. Second, ignoring the mutual coupling among RIS elements in the optimization process leads to performance degradation, particularly when the inter-element spacing of RIS is small. Third, the unilateral approximation is highly accurate even when the distance between the transmitter and RIS is only a few wavelengths. Therefore, it can be employed in practice to simplify the channel model without loss of accuracy.
@@ VII Conclusion
In this paper, we studied the architecture design and optimization of BD-RIS under the general physics-consistent model derived in [28], which captures practical non-linearities such as imperfect matching and mutual coupling. To this end, we derived a compact reformulation of the model that makes the impact of RIS explicit, thereby facilitating both optimization and analysis. Based on this formulation, we make the following two key contributions. First, we prove that band-connected RIS is optimal under the general physics-consistent model, which unifies and generalizes existing results in [12, 17, 14]. Second, we develop optimization algorithms for different systems under the general physics-consistent model. This enables us to give a comprehensive evaluation of the impact of various effects and approximations. In particular, our simulations demonstrate that the unilateral approximation widely adopted in the literature remains highly accurate even when the transmission distance is only a few wavelengths. In contrast, mutual coupling among RIS elements has a significant influence on system performance: stronger mutual coupling can enhance performance under Rayleigh fading channels, and neglecting it during optimization leads to performance degradation.

### v2

@@ V-B Optimization of Single-Stream MIMO Systems
For the general physics-consistent channel model given by Proposition 1, $\bar{\mathbf{H}}_{RT}$ has a more complicated expression and is generally nonzero. As a result, closed-form solutions are generally unavailable. Nevertheless, we develop in the following a novel algorithm that is able to attain the global optimal solution of (21) (for fully- and tree-connected RIS). This is highly nontrivial due to the nonconvex nature of the problem, in particular the presence of complicated constraints and the multiplicative coupling of the variables in the objective function. As discussed earlier, the existing approach in [7] cannot provide such a global optimality guarantee.
Specifically, we first eliminate variable $\mathbf{g}$ by substituting the optimal $\mathbf{g}$ into the objective function of (21). By the Cauchy–Schwarz inequality, the objective function of (21) is maximized with respect to $\mathbf{g}$ when $\mathbf{g}$ is aligned with $\left(\bar{\mathbf{H}}_{RT}+\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}\right)\mathbf{w}$, i.e.,
$\mathbf{g}^{*}(\bm{\Theta},\mathbf{w})=\frac{\left(\bar{\mathbf{H}}_{RT}+\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}\right)\mathbf{w}}{\|\left(\bar{\mathbf{H}}_{RT}+\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}\right)\mathbf{w}\|_{2}}.$
Substituting $\mathbf{g}^{*}(\bm{\Theta},\mathbf{w})$ into the objective function of (21) yields the following equivalent problem:
@@
The constant $P_{T}$ does not affect the optimal solution and is thus omitted for brevity. To further handle the multiplicative coupling between the variables $\bar{\bm{\Theta}}$ and $\mathbf{w}$ in the objective function, we introduce an auxiliary variable $\mathbf{u}=\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}\mathbf{w}$, then problem (22) transforms to
@@
where the non-convex rank-one constraint $\text{rank}(\mathbf{X})=1$ is relaxed. A well-known result for a complex SDP of the form (26) is that, there exists an optimal solution $\mathbf{X}^{\star}$ with $\operatorname{rank}(\mathbf{X}^{\star})\leq\sqrt{m}$ (as long as the problem is feasible), where $m$ denotes the number of linear constraints [33, 34]. Since $m=2$ in (26), there exists a rank-one optimal solution, and thus the SDR is tight. By decomposing $\mathbf{X}^{*}=\mathbf{x}^{*}{\mathbf{x}^{*}}^{H}$, we get the optimal solution to (24).
@@
Solving the ($N_{T}+N_{I}$)-dimensional semidefinite program (SDP) in (26) is computationally expensive, as the number of RIS elements is typically large. This remark will show that, by carefully exploiting the problem structure, it suffices to solve an SDP with a much lower dimension of $N_{T}+N_{R}$, where $N_{R}\ll N_{I}$ in practice.
@@ V-C Optimization of Multiuser MIMO Systems
Finally, we discuss sum-rate maximization for multiuser MIMO systems. For ease of presentation, we assume that each user has a single antenna. In this case, $N_{R}$ is the number of users in the system. The sum-rate maximization problem can be formulated as
[…]
$\|\mathbf{W}\|_{F}^{2}\leq P_{T},$
where $\mathbf{h}_{k}(\bar{\bm{\Theta}})^{H}$ is the $k$-th row of $\mathbf{H}(\bar{\bm{\Theta}})=\bar{\mathbf{H}}_{RT}+\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}}\bar{\mathbf{H}}_{IT}$.
Various algorithms have been proposed to solve the sum-rate maximization problem under the conventional channel model. The new challenge introduced by the general physics-consistent model lies in that the virtual scattering matrix $\bar{\bm{\Theta}}$ involved in the channel model is related to the actual susceptance matrix $\mathbf{B}_{I}$ through a more intricate relationship, given by (13) and (14). For fully-connected RIS, this constraint reduces to the unitary and symmetric constraints on $\bar{\bm{\Theta}}$, and thus existing approaches remain directly applicable. However, for more general architectures, (13) and (14) must be explicitly taken into account. In particular, for group-connected RIS, the block-diagonal structure of the susceptance matrix $\mathbf{B}_{I}$ does not generally translate into a block-diagonal structure of $\bar{\bm{\Theta}}$, making existing approaches that rely on such structural properties not directly applicable.
Among the existing methods, our recent work in [10] proposed an ADMM-based optimization framework that is applicable to arbitrary RIS architectures, which was shown to achieve a favorable trade-off between computational complexity and performance. In the following, we generalize this algorithmic framework to accommodate the general physics-consistent channel model, by carefully dealing with the coupling between $\bar{\mathbf{B}}_{I}$ and $\mathbf{B}_{I}$ in (14).
@@
As in [10], we introduce an auxiliary variable $\mathbf{U}=(\bar{\mathbf{H}}_{RI}\bar{\bm{\Theta}})^{H}$ to reformulate (30) as
@@
where $\mathbf{h}_{k}(\mathbf{U})^{H}$ is the $k$-th row of $\mathbf{H}(\mathbf{U})=\bar{\mathbf{H}}_{RT}+\mathbf{U}^{H}\bar{\mathbf{H}}_{IT}.$ To solve problem (31), we employ the ADMM-based algorithmic framework in [10, Eqs. (11a) – (11f)], which first simplifies the sum-rate objective function using fractional programming technique [35, 36] and penalizes the bilinear constraint (31b) into the objective function via the augmented Lagrangian (AL) function, and then alternately updates the variables until convergence. The only difference lies in the update of $(\bar{\mathbf{B}}_{I},\mathbf{B}_{I})$, where the architecture-specific constraint on $\mathbf{B}_{I}$, i.e., $\mathbf{B}_{I}\in\mathcal{B}$, is coupled with $\bar{\mathbf{B}}_{I}$ through the linear constraint (14). In the following, we focus on the update of $(\bar{\mathbf{B}}_{I},\mathbf{B}_{I})$ involved in the ADMM algorithm. Details on the updates for the other variables are omitted for brevity and can be found in [10].
@@
The adopted framework can be readily extended to the multiuser MIMO scenario by applying the fractional programming technique in its matrix form [37]. This generalization has been discussed in [10, Section V-B].
@@ VI Simulation
In this section, we present simulation results to examine the impact of different EM effects and approximations on system performance, using the algorithms proposed in Section V. We focus on fully-connected RIS and evaluate its performance under different channel models. In particular, we investigate two key effects: mutual coupling (MC) among RIS elements and unilateral approximation (UA). To this end, we assume that the source impedances at the transmitter and load impedances at the receiver are equal to the reference impedance $Z_{0}$, set as $Z_{0}=50\,\Omega$, which gives $\mathbf{Z}_{R}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{T}=Z_{0}\mathbf{I}$, and the transmit and receive antennas are perfectly matched with no mutual coupling, which gives $\mathbf{Z}_{TT}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{RR}=Z_{0}\mathbf{I}.$ In addition, the direct link between the transmitter and receiver is assumed to be fully obstructed, i.e., $\mathbf{Z}_{RT}=\mathbf{0}$. We focus on the following three channel models:
@@
Throughout the simulations, the transmit power is set to $P_{T}=20$ dBm and the noise power is $\sigma^{2}=-80$ dBm. To thoroughly analyze the impact of mutual coupling and unilateral approximation, we consider two scenarios with different modeling assumptions for the channels $\mathbf{Z}_{RI}$ and $\mathbf{Z}_{IT}$.
  1. 1.
@@
    First, we consider a common scenario in which the distances between the transmitter, RIS, and receiver are sufficiently large. In this case, the elements of $\mathbf{Z}_{RI}$ and $\mathbf{Z}_{IT}$ are generated as independent Gaussian random variables [17], i.e., $\mathbf{Z}_{RI}]_{i,j}\sim\mathcal{CN}(0,\rho_{RI}),$ and $\mathbf{Z}_{IT}]_{i,j}\sim\mathcal{CN}(0,\rho_{IT})$, with path gain $\rho_{IT}=4Z_{0}^{2}10^{-8}$ and $\rho_{RI}=4Z_{0}^{2}10^{-4}$. The corresponding results are given in Figs. 3 – 5.
  2. 2.
    Then, we consider a scenario in which the distance between the transmitter and the RIS is on the order of the wavelength. The receiver is still assumed to be located far from the RIS. Specifically, we set the transmitter parallel to the $y$-axis, centered at $(d,0,r\lambda)$, where $d$ is the inter-element spacing of RIS, $\lambda$ is the wavelength, and $r$ is a factor determining the distances between transmitter and RIS. The matrix $\mathbf{Z}_{IT}$ is then computed according to the model in (37), and all other components of $\mathbf{Z}$ are generated as in 1). The results are given in Fig. 6.
To investigate the performance of different channel models for different systems, we solve (20), (21), and (30) under both the general physics-consistent model $\mathbf{H}$ and the approximate models $\mathbf{H}_{\text{app},2}$ and $\mathbf{H}_{\text{app},3}$. Specifically, the approximate solutions are obtained by applying the algorithms in Section V to solve the receive power or sum-rate maximization problems using the approximate channel models $\mathbf{H}_{\text{app,2}}$ and $\mathbf{H}_{\text{app,3}}$, i.e., $\bar{\mathbf{H}}_{RT}$, $\bar{\mathbf{H}}_{RI}$, $\bar{\mathbf{H}}_{IT}$, and $\bar{\bm{\Theta}}$ in problems (20), (21), and (30) are replaced by their approximate counterparts.
@@
In Fig. 3, we depict the receive power for SISO and single-stream MIMO systems and the sum rate for multiuser MISO systems achieved by the three channel models. To better investigate the impact of mutual coupling among RIS antennas, we consider two different values of inter-element spacing, $d=\lambda/2$ and $d=\lambda/4$, respectively. As shown in the figure, the performance with and without the unilateral approximation is almost indistinguishable in all cases. In contrast, mutual coupling among RIS antennas has a noticeable impact on the performance, particularly when the inter-element spacing is small. The presence of mutual coupling improves system performance compared with the case without coupling, and stronger coupling leads to more significant performance gains. We remark here that the benefits of mutual coupling under Rayleigh fading channels were previously reported for SISO systems in [17]. Our results demonstrate that the same conclusion also holds for MIMO and multiuser scenarios.
@@
In Fig. 4, we further evaluate the quality of the approximate solutions obtained using the channel models $\mathbf{H}_{\text{app,2}}$ and $\mathbf{H}_{\text{app,3}}$ to better visualize the accuracy of different approximations. The quality of each approximate solution is quantified by applying it to the general physics-consistent channel model $\mathbf{H}$ and computing its relative performance compared with the solution obtained from $\mathbf{H}$. From Fig. 4, we can draw similar conclusions as those from Fig. 3. First, the unilateral approximation has a negligible effect in all cases, yielding nearly 100% relative performance compared to the solution of the general physics-consistent model. Second, ignoring mutual coupling among RIS elements leads to performance degradation, particularly when the inter-element spacing is small. For example, when $d=\lambda/4$, the solution from the mutual-coupling-unaware model achieves only around $65\%$ of the performance.
@@
In all the above figures, the distances between transmitter, RIS, and receiver are set to be much larger than the wavelength, which is typically the case in practice. It is commonly accepted that the unilateral approximation accurately represents real world channels under this condition, as also validated by our results. In Fig. 6, we further investigate the range of validity of the unilateral approximation by considering an extreme scenario where the distance between the transmitter and RIS is on the order of the wavelength. We assume that mutual coupling is aware and focus on the effect of unilateral approximation by comparing the performance achieved with models $\mathbf{H}_{\text{app,2}}$ and $\mathbf{H}$. The number of RIS elements is fixed as $N_{I}=64$.
As shown in Fig. 6, the unilateral approximation remains accurate across a broad range, even when the distance between the transmitter and RIS is on the order of the wavelength. For example, even at $0.1\lambda$, the solution obtained under the unilateral approximation still achieves over 95% of the performance of the exact solution for the SISO system. Although the accuracy decreases with the the number of transmit/receive antennas and users, it is still sufficiently high for all normal RIS-aided wireless communication systems.
We remark that we have also investigated the impact of unilateral approximation and mutual coupling under different RIS architectures (other than the fully-connected RIS). The results exhibit trends consistent with those in Figs. 46. This is expected, since the unilateral approximation and mutual coupling are related to EM effects among/of the transmit, RIS, and receive antennas, and are irrelevant to the circuit topology of the reconfigurable impedance network. For brevity and clarity, we omit these additional figures.
In summary, we can draw the following conclusions from the simulation results regarding the effects of mutual coupling and unilateral approximation. First, stronger mutual coupling among RIS elements can enhance system performance for all SISO, MIMO, and multiuser MISO systems under Rayleigh fading channels. Second, ignoring the mutual coupling among RIS elements in the optimization process leads to performance degradation, particularly when the inter-element spacing of RIS is small. Third, the unilateral approximation is highly accurate even when the distance between the transmitter and RIS is only a few wavelengths. Therefore, it can be reliably employed in normal RIS-aided wireless communication systems to simplify the channel model without loss of accuracy.
@@ VII Conclusion
In this paper, we studied the architecture design and optimization of RIS under the general physics-consistent model derived in [29], which captures practical non-linearities such as imperfect matching and mutual coupling. To this end, we derived an explicit expression of the model that makes the impact of the RIS explicitly visible. Under the common assumptions that the RIS is lossless and reciprocal, we further derived a compact reformulation of the model, thereby facilitating both optimization and analysis. Based on this formulation, we make the following two key contributions. First, we prove that band-connected RIS is optimal under the general physics-consistent model, which unifies and generalizes existing results in [12, 17, 14]. Second, we develop optimization algorithms for different systems under the general physics-consistent model. This enables us to give a comprehensive evaluation of the impact of various effects and approximations. Our simulations demonstrate that the unilateral approximation widely adopted in the literature is highly accurate for normal RIS-aided wireless communication systems. In contrast, mutual coupling among RIS elements has a significant influence on system performance: stronger mutual coupling can enhance performance under Rayleigh fading channels, and neglecting it during optimization leads to performance degradation.